### v1

# Human Motion Capture from Loose and Sparse Inertial Sensors with Garment-aware Diffusion Models

Abstract
Motion capture using sparse inertial sensors has shown great promise due to its portability and lack of occlusion issues compared to camera-based tracking. Existing approaches typically assume that IMU sensors are tightly attached to the human body. However, this assumption often does not hold in real-world scenarios. In this paper, we present a new task of full-body human pose estimation using sparse, loosely attached IMU sensors. To solve this task, we simulate IMU recordings from an existing garment-aware human motion dataset. We developed transformer-based diffusion models to synthesize loose IMU data and estimate human poses based on this challenging loose IMU data. In addition, we show that incorporating garment-related parameters while training the model on simulated loose data effectively maintains expressiveness and enhances the ability to capture variations introduced by looser or tighter garments. Experiments show that our proposed diffusion methods trained on simulated and synthetic data outperformed the state-of-the-art methods quantitatively and qualitatively, opening up a promising direction for future research.
1 Introduction
Human pose estimation is essential in healthcare, sports, ergonomics, and AR/VR applications. Traditionally, it has been achieved via images or videos  Lan et al. (2023); Sosa and Hogg (2023); Pavllo et al. (2019); Wang et al. (2020); Zhang et al. (2022a); Artacho and Savakis (2020); Chen et al. (2021); Liu et al. (2021), AR/VR devices Zheng et al. (2023); Jiang et al. (2022a); Dai et al. (2024); Millerdurai et al. (2024) and Inertial Measurement Units (IMUs). In recent years, various models have relied on IMUs as a portable and privacy-friendly solution Yi et al. (2024); De Marchi et al. (2024). Some notable examples like PIP Yi et al. (2022), TIP Jiang et al. (2022b), TransPose Yi et al. (2021), and DynaIP Zhang et al. (2024) focus on full-body pose estimation from a minimal number of IMUs. However, they all share a common limitation: the reliance on IMUs that are tightly strapped to the body, thus sacrificing comfort. In contrast, IMUs loosely attached to garments allow individuals to perform everyday activities while wearing regular clothing without feeling restricted or encumbered. In this way, the tracking of human motion in diverse environments, beyond controlled settings, could be facilitated.
Figure 1: We present a novel task of full-body human pose estimation using sparse IMUs that are loosely attached to garments. To address this challenge, we propose methods for simulating and synthesizing loose IMU signals using diffusion models. Leveraging both simulated and synthetic data, we design diffusion models conditioned on IMU data alone or IMU data and clothing model parameters, enabling robust human motion estimation. We evaluate the effectiveness of our approach across three datasets: a simulated dataset, a real-world upper-body dataset, and a real-world full-body dataset.
Loose Inertial Poser (LIP)  Zuo et al. (2024) explores this setup, focusing on upper-body pose using four IMUs. The Secondary Motion Autoencoder (Semo-AE) they proposed tackles the lack of sufficient real-world training data, as simulated data often fails to capture the full variability seen in actual human movements. However, this model was trained exclusively on recordings from tightly and loosely attached IMUs, disregarding other factors that influence secondary motion, such as the individual’s pose or garment-related characteristics. In addition, it does not address full-body tracking.
To overcome these limitations, we propose a real-time method using conditional diffusion models for full- and upper-body pose estimation from loosely worn IMUs. We train our models on simulated, synthetic, and real datasets, using a secondary diffusion model to generate realistic loose-wear data. While moderately loose conditions represented by simulated and synthetic data provide satisfactory performance, training the model on real-world data reveals a problem with limited motion dynamics and convergence of the model towards the mean prediction. This motivated us to shift from the goal of generating realistic-looking synthetic data to incorporating garment characteristics into the model training. We evaluate our models on simulated Mahmood et al. (2019), real-world upper-body Zuo et al. (2024), and real-world whole-body Lorenz et al. (2022) datasets, demonstrating improved generalization and robustness to varying garment looseness.
To summarize, our contributions are as follows:
  1. 1.
    the new task of full-body pose estimation using sparse IMU sensors that are loosely attached. Previous methods assumed tightly attached IMUs or only attempted partial reconstruction, whereas our work tackles a more challenging and realistic scenario.
  2. 2.
    a training strategy that applies three distinct types of loosely attached IMU data: simulated data derived from existing garment-aware human motion datasets, synthetic data generated by our secondary diffusion model, and real-world recordings.
  3. 3.
    a transformer-based diffusion model that estimates full-body poses from loosely attached IMUs. We show that incorporating garment-related parameters significantly enhances pose accuracy, even amid variations in signal noise and differences in garment fit across individuals.
  4. 4.
    several experimental evaluations that show that our diffusion models outperform the state-of-the-art (SOTA) methods. We also provide comprehensive ablation studies to thoroughly analyze the effectiveness of our methods.
2 Related Work
Human pose estimation from wearable sensors.
Models like DIP, TransPose, TIP, PIP, and DynaIP focus on real-time full-body pose reconstruction using six IMUs. Researchers have noted that this task is under-constrained Huang et al. (2018); Mollyn et al. (2023), as many poses can match the same IMU readings. The scarcity of ground-truth data is often addressed by leveraging simulated IMU data derived from mocap datasets. Additionally, modeling temporal dependencies can sometimes impede real-time predictions. TransPose Yi et al. (2021) highlights signal sparsity and noise, while TIP Jiang et al. (2022b) introduces Transformers to improve temporal consistency and reduce drift. PIP Yi et al. (2022) is the first real-time, physics-aware model estimating both pose and forces. DynaIP Zhang et al. (2024) shifts from synthetic to real IMU data and minimizes the influence of less relevant joints to reduce ambiguity. Beyond standalone IMUs, recent efforts include VR wearables Jiang et al. (2022a, 2024b, 2024a), camera-based systems Hollidt et al. (2024), and hybrid setups combining IMUs with UWB sensors Armani et al. (2024); Armani and Holz (2024).
Pose estimation from loosely attached IMUs.
To the best of our knowledge, the only method tailored for loosely worn IMUs is Loose Inertial Poser (LIP) Zuo et al. (2024). It introduces a Secondary Motion Autoencoder that treats secondary motion as Gaussian noise, using latent space learning to model variation across body shapes and garment types. A temporal coherence mechanism ensures smooth, realistic motion. Trained on AMASS and evaluated on real upper-body data, LIP achieved $<20$° joint rotation error and 10.6 cm position error and outperformed several SOTA models (PIP, TIP, DIP, TransPose) trained on simulated loose IMU data. LIP uses LSTMs, which suit sequential IMU inputs. In contrast, our pipeline leverages modern generative models to better capture uncertainty and complex motion patterns.
Diffusion models for human motions.
Diffusion models have recently shown promise for motion generation tasks Zhang et al. (2022b); Tevet et al. (2022); Tseng et al. (2022); Yuan et al. (2023); Zhou et al. (2025); Kim et al. (2023). In pose estimation, they have been applied to HMD-based setups Feng et al. (2024); Du et al. (2023), where sparse upper-body tracking must be used to infer full-body motion. These models perform well under such under-constrained conditions. DiffusionPoser Wouwe et al. (2024) is the most relevant work here—it uses IMUs and pressure insoles with an unconditional diffusion model that supports flexible sensor configurations and is robust to signal degradation. Unlike DiffusionPoser, which uses tightly worn sensors, our method employs conditional diffusion models trained on both generated and real-world data with loosely attached IMUs.
3 Proposed Method
3.1 Method Overview
We estimate full-body human pose from loosely attached garment-mounted IMUs using acceleration and orientation inputs to predict SMPL joint rotations.
Training data is generated by simulating IMU signals from a garment-aware motion capture dataset. Our approach uses two conditional diffusion models: a primary model for pose estimation and a secondary model for generating realistic loose-wear IMU signals. The secondary model, conditioned on tight-wear IMU data and pose, synthesizes loose-wear signals. These are used to train the primary model, which predicts pose, tight-wear IMU signals, and joint positions. A garment-aware variant further conditions on clothing parameters to enhance realism.
3.2 Diffusion Framework
We adopt a Denoising Diffusion Probabilistic Model (DDPM) Ho et al. (2020) as the generative backbone. This model operates by progressively corrupting data through a forward process, where Gaussian noise is incrementally added to the data over successive timesteps, forming a series of latent variables $z_{t}$. Starting from a clean data sample $x_{0}$, the forward diffusion process follows a Markov chain that applies Gaussian noise at each timestep $t$, defined as:
$q(z_{t}|x_{0})=\mathcal{N}\left(z_{t};\sqrt{\alpha_{t}}x_{0},(1-\alpha_{t})\mathbf{I}\right),$
where $\alpha_{t}$ is a noise scheduling coefficient that controls the amount of noise added at each step, and $t\in[0,T]$ denotes the progression of time. As $t$ approaches the maximum timestep $T$, the sample $z_{T}$ approximates a Gaussian distribution $\mathcal{N}(0,\mathbf{I})$, effectively turning the original data into pure noise.
To recover the original data, a reverse process is learned, which denoises the latents step by step. The reverse diffusion process is modeled as a sequence of conditional probabilities $p_{\theta}(z_{t-1}|z_{t})$, where $\theta$ represents the parameters learned by the model. This process is parameterized by a neural network, which predicts either the noise-free data $x_{0}$ or the noise component $\epsilon_{t}$ added during the forward process. For the pose estimation task, we opted to predict the signal itself $x_{0}$, as this facilitates a more straightforward application of the objective function.
Figure 2 shows the architecture of the neural network used in all the proposed diffusion models. This is a transformer-based autoencoder with an arbitrary number of transformer blocks in the encoder and decoder networks.
The embedding layer consists of four convolutional layers applied to different parts of the input sequence and the condition, for efficient feature extraction. These parts refer to distinct signals concatenated before being fed into the diffusion model. The use of causal self-attention and residual connections enables effective sequence modeling. To preserve the sequence information, positional encoding is consistently applied before passing the final vector to the first transformer block of the encoder.
Figure 2: Denoiser architecture used in all diffusion models.
3.3 Data Simulation and Generation
We simulate both tightly and loosely attached IMU sensor data for poses from the AMASS dataset. The IMU observations include the acceleration and rotation information from 6 sensors:
$\text{IMU}(t)=\left(\{a_{1}(t),a_{2}(t),\ldots,a_{6}(t)\},\{R_{1}(t),R_{2}(t),\ldots,R_{6}(t)\}\right).$
The sensors are positioned on the left and right forearm, sternum, pelvis, and left and right lower leg. For the simulation of clothing effects, we employ TailorNet’s Patel et al. (2020) Male Shirt and Male Pants models to maintain consistency with the LIP paper. Specifically, we select the SMPL body model Loper et al. (2015) and utilize the TallThin physique and garment style with zero $\gamma$ parameter. Following LIP’s loose data simulation, we select four neighboring vertices from the clothing model to define the IMU orientation through cross-products, while the geometric center of these vertices serves as its position. Similarly, we obtain tight IMU data directly specifying the corresponding vertex on the AMASS body mesh.
Body Part Simulated Data [] Real-World Data []
Left Forearm 21.451 40.890
Right Forearm 21.575 49.545
Sternum 21.094 27.395
Pelvis 12.195 25.224
Left Lower Leg 15.819 35.449
Right Lower Leg 13.512 36.002
Table 1: Orientation difference between tightly and loosely worn IMU sensors’ rotations, calculated as the angular difference in degrees. The difference is computed by converting the rotations to rotation matrices, finding the offset using matrix multiplication, and then applying the axis-angle representation to determine the angle.
The orientation difference between the simulated tight and loose data is significantly smaller than that found in real-world data provided by Lorenz et al. Lorenz et al. (2022), as shown in Table 1. This discrepancy effectively represents tighter-fitting garments, making the simulation less reflective of real-world conditions. To align our dataset better with real-world scenarios, we developed a diffusion model (Figure 3) trained on real-world data provided by Lorenz et al. Lorenz et al. (2022). This model generates loose-wear IMU data $c_{l}$ conditioned on tight-wear IMU data $c_{t}$ and pose information $\theta$ represented as a quaternion. While the LIP model generates synthetic loose data by introducing noise into the tight data, we believe the pose provides essential secondary motion context. We included pose as the condition, as it directly influences how garments move and interact with the body. The objective function is the simple L1 reconstruction loss:
$\mathcal{L}_{L1}=\|c_{l}-\hat{c}_{l}\|_{1}=\sum_{i=1}^{N}|c_{l}^{i}-\hat{c}_{l}^{i}|.$ (1)
Figure 3: Data generation scheme.
Despite the improved offset, we observed that the generated data lacked sufficient diversity, often reflecting scenarios with very wide garment configurations. Therefore, we introduced a parameter $\alpha\sim\text{Uniform}(0,1)$, which modulates the contribution of the simulated and generated data. By varying $\alpha$, we can adjust the tightness or looseness of the clothing representation, effectively blending more or less of the simulated or generated signal. The final training data is obtained using the following formula:
$c_{i}=\alpha c_{s}+(1-\alpha)c_{l},$ (2)
where $c_{i}$ represents the interpolated synthetic loose data, $c_{s}$ is the simulated loose data, and $c_{l}$ is the loose data generated by our diffusion model.
3.4 Conditional Diffusion Models for Pose Estimation
To estimate the pose based on the observations from loosely attached sensors, we developed a conditional diffusion model (Figure 4). These sensor observations serve as conditions throughout the diffusion process, enabling the model to generate not only pose estimates but also predictions of tight IMU data and joint positions. Given the conditional nature of this model, where predictions fully rely on the provided input, a flexible number of available sensors cannot be accommodated. Therefore, we propose two distinct models—one for the upper-body pose estimation and another for the full-body pose estimation.
The upper-body model predicts the positions and rotations of 14 upper-body joints, including leaf joints, while the whole-body model predicts all 24 joints, as defined by the SMPL body model Loper et al. (2015). The loss function for the upper-body model is defined as
$\begin{split}\mathcal{L}=\alpha_{r_{r}}\mathcal{L}_{r_{r}}+\alpha_{r_{j}}\mathcal{L}_{r_{j}}+\alpha_{p_{a}}\mathcal{L}_{p_{a}}\\ +\alpha_{p_{j}}\mathcal{L}_{p_{j}}+\alpha_{t}\mathcal{L}_{t}+\alpha_{c}\mathcal{L}_{c}.\end{split}$ (3)
Here, $\mathcal{L}_{r_{r}}$ and $\mathcal{L}_{r_{j}}$ represent the reconstruction errors of the root joint rotation and the remaining joints’ rotations, respectively. The terms $\mathcal{L}_{p_{a}}$ and $\mathcal{L}_{p_{j}}$ capture the position errors for the arms and all other joints, respectively, while $\mathcal{L}_{t}$ corresponds to the reconstruction error of the tightly attached IMU observations. Finally, $\mathcal{L}_{c}$ denotes the consistency regularization, calculated as the difference between the model’s predictions when conditioned on the original sensor data and a noisy version of the data, where Gaussian noise scaled by 0.3 is added. The coefficients are set as follows: $\alpha_{r_{r}}=2$, $\alpha_{r_{j}}=1$, $\alpha_{p_{a}}=2$, $\alpha_{p_{j}}=1$, $\alpha_{t}=1$, and $\alpha_{c}=3$.
The loss function for the whole-body model follows the same structure, with $\mathcal{L}_{p_{al}}$ (position error of arms and legs) replacing $\mathcal{L}_{p_{a}}$ (position error of arms) to account for the position error of all extremities. Similarly, $\mathcal{L}_{r_{j}}$ now incorporates the rotation estimation errors for the lower body as well. All loss functions are based on L1 loss. Additionally, we assign higher weights to the root joint’s rotation and the positions of the extremities, as these proved to be particularly challenging for accurate estimation.
Figure 4: Conditional diffusion model.
3.5 Garment-Aware Conditional Model
The conditional model trained on real-world whole-body IMU data revealed issues with very loose garment configurations, particularly with the root joint sensor, which is placed on top of the jacket and over the pants, leading to highly noisy recordings. Similarly, sensors placed on the extremities often provide misleading observations due to garment movement. To mitigate this intrinsic uncertainty, we try to enhance the conditional model by integrating garment-related information. To achieve this, we utilize seven simulations generated using TailorNet’s MaleShirt model. Six of these simulations correspond to a TallThin physique, denoted as [180, 22], where 180 represents height in centimeters and 22 refers to the body mass index (BMI). These simulations encompass six garment style variations, with $\gamma$ values of 0, 5, 10, 15, 20, and 24. The final simulation represents a ShortFat physique, denoted as [160, 30], with a $\gamma$ value of 0. We focus on four upper-body IMU sensors for these simulations.
The training scheme remains consistent with that shown in Figure 4, except for the condition dimension, which now expands to 39 (comprising 36 loose sensor values, $\gamma$, height, and BMI). As a result, the training dataset is expanded by a factor of seven, yielding approximately 47 million samples.
3.6 Real-Time Inference
We aim to achieve real-time inference and seamless predictions, which is particularly challenging when working with diffusion models. On one hand, these models require time to denoise the input; on the other hand, they do not keep track of previous estimations. As a result, motion predictions are disconnected, leading to significant jitter. For all these reasons, we opted for the frame-by-frame prediction strategy. We use a sliding window of length $N$ (equal to the training sequence length) and shift it by one frame at a time, using the last frame’s prediction as the estimate for the current timestep. To enhance temporal consistency and reduce jitter, we incorporate a masking strategy during inference, as proposed by Van Wouwe et al. Wouwe et al. (2024), which suggests fixing past predictions to produce smoother, more coherent movements.
4 Experiments
4.1 Datasets
AMASS dataset.
AMASS is a large-scale motion capture dataset with more than 40 hours of diverse motions. Specifically, we use the 10 subsets (BMLrub, CMU, EyesJapan, HumanEva, MPIHDM05, MPIMoSh, SFU, SSM, Transitions_mocap, BMLmovi) for training, and the rest 5 subsets TotalCapture, ACCAD, TCD_handMocap, BMLhandball, and MPI_Limits, for the evaluation. These five datasets ensure diversity in motion capture, ranging from everyday activities in ACCAD to extreme poses in MPI_Limits, from detailed hand motions in TotalCapture and TCD_handMocap to dynamic sports data in BMLhandball.
LIP dataset.
The LIP dataset, provided by the authors of the LIP paper Zuo et al. (2024), contains IMU signals recorded from four sensors placed on the upper body: left forearm, right forearm, back, and waist. The dataset includes two wearing styles: zipped and unzipped, and participants performed walking, running, jumping, boxing and ping-pong, along with five sessions of free-form movement. In total, the dataset comprises 212,496 frames recorded at 30 frames per second.
Whole-body dataset.
Since the LIP dataset is limited to upper-body motions, we utilized a whole-body dataset provided by Lorenz et al. Lorenz et al. (2022), with 12 IMUs placed on loose-fitting clothing to capture whole-body movements. Ten participants performed various upper and lower body movements and a 15-minute activity sequence. The pelvis sensor was attached to the jacket, not the trousers, and the work suit sleeves were unbuttoned. The dataset includes 1,137,418 frames recorded at 60 fps. We utilize only the rotation data from this dataset, since global acceleration measurements are not available. To obtain the SMPL parameters, we apply inverse kinematics to the global joint orientations provided in the dataset and then map with the corresponding joint names.
4.2 Evaluation Metrics
Following the evaluation protocol of LIP Zuo et al. (2024), we compute the angular error, which is defined as the Mean Per Joint Rotation Error (MPJRE) in degrees. We also calculate the position error as the Mean Per Joint Position Error (MPJPE) in centimeters. To measure the smoothness of the predicted motions, we include two additional metrics: the Mean Per Joint Velocity Error (MPJVE) in cm/s, and Jitter, which quantifies the changes in acceleration over time, expressed in $10^{2}\,\text{m/s}^{3}$. For the upper-body evaluation, we average results across 11 joints, while the whole-body evaluation considers all 24 SMPL model joints.
4.3 Evaluation Results
In this section, we compare our proposed methods and the state-of-the-art pose estimation methods using loosely attached IMUs. Note that LIP has previously demonstrated superior performance over methods designed for tightly-attached IMUs Mollyn et al. (2023); Yi et al. (2022); Huang et al. (2018); Yi et al. (2021); Jiang et al. (2022b).
Table 2 presents the evaluation metrics for the best-performing models across all three datasets. The terms Sim, Syn, and Real in our model names indicate whether the model was trained on simulated, synthetic, or real loose data.
Table 3 compares models trained and tested on simulated AMASS data. We evaluate methods designed for loosely attached sensors (Ours, LIP) against those developed for tight-fitting setups (IMUPoser Mollyn et al. (2023), PIP Yi et al. (2022)) under varying levels of sensor dropout.
Figures 5, 6 and 7 illustrate the qualitative performance of our conditional models, including the upper-body and whole-body garment-unaware models, as well as the upper-body garment-aware model.
Figure 5: Qualitative performance of the upper-body models on all three evaluation datasets.
Figure 6: Qualitative performance of the whole-body models on the simulated and real-world datasets.
Figure 7: Qualitative performance of the garment-aware conditional upper-body model on the real-world datasets.
     Datasets      Metrics      Upper-body Sim (Ours)      Upper-body Syn (Ours)      Upper-body LIP      Whole-body Sim (Ours)      Whole-body Syn (Ours)      Whole-body LIP      Upper-body Real (Ours)      Whole-body Real (Ours)
AMASS
MPJRE
MPJPE
MPJVE
Jitter
GT Jitter
14.46 ± 6.00
9.58 ± 6.33
26.66
63.03
137.07
15.90 ± 6.75
12.12 ± 9.50
28.18
54.65
137.07
15.05 ± 5.81
10.58 ± 6.46
27.65
95.83
137.07
13.65 ± 4.47
10.36 ± 5.32
27.54
55.76
140.76
14.61 ± 4.72
12.54 ± 6.38
30.64
47.00
140.76
13.67 ± 4.47
12.78 ± 8.76
31.40
93.37
140.76
27.52 ± 6.89
15.40 ± 9.57
31.35
44.02
137.07
23.28 ± 4.94
18.16 ± 9.32
32.89
34.34
140.76
LIP
MPJRE
MPJPE
MPJVE
Jitter
GT Jitter
20.68 ± 5.66
12.17 ± 6.01
32.88
112.51
33.41
19.93 ± 5.68
10.65 ± 5.86
32.68
88.40
33.41
20.85 ± 6.36
11.49 ± 6.70
41.99
233.40
33.41
__
__
__
__
__
__
__
__
__
__
__
__
__
__
__
28.31 ± 5.99
16.50 ± 7.42
35.26
61.89
33.41
__
__
__
__
__
Whole Body
MPJRE
MPJPE
MPJVE
Jitter
GT Jitter
31.21 ± 7.52
28.80 ± 16.02
22.68
23.61
43.97
30.27 ± 8.21
25.21 ± 14.10
21.12
15.06
43.97
33.18 ± 8.84
26.39 ± 15.59
23.97
23.03
43.97
24.72 ± 5.70
28.65 ± 14.47
26.97
34.80
54.70
25.16 ± 5.95
28.60 ± 14.37
28.10
30.32
54.70
25.97 ± 7.02
26.23 ± 14.5
30.48
50.59
54.70
__
__
__
__
__
__
__
__
__
__
Table 2: Numerical comparisons between our methods and the state-of-the-art method LIP. We show results for the upper-body as well as the whole-body settings. We also provide the performance of the models trained only on the real-world data to show the benefit of our data simulation and synthesis strategies.
Number of missing sensors      Metrics AMASS Whole-body real-world dataset
Whole-body Sim (Ours) Whole-body LIP IMUPoser PIP IMUPoser PIP
0 MPJRE MPJPE 13.65 ± 4.47 10.36 ± 5.32 13.67 ± 4.47 12.78 ± 8.76 12.08 ± 4.77 7.37 ± 4.43 17.38 ± 5.03 9.62 ± 4.84 38.86 ± 5.30 44.03 ± 8.77 43.43 ± 7.27 44.62 ± 9.53
1 MPJRE MPJPE 14.02 ± 4.58 10.71 ± 5.47 15.41 ± 4.86 14.78 ± 8.84 12.11 ± 4.86 7.54 ± 4.49 20.47 ± 4.35 23.84 ± 8.51
2 MPJRE MPJPE 14.56 ± 4.47 11.76 ± 5.49 17.14 ± 4.89 17.26 ± 9.06 12.75 ± 5.04 8.52 ± 4.76 24.84 ± 4.39 37.82 ± 9.91
3 MPJRE MPJPE 17.13 ± 4.34 16.48 ± 6.58 19.71 ± 4.62 21.31 ± 9.19 16.05 ± 5.99 12.26 ± 5.81 30.42 ± 4.56 47.14 ± 9.82
Table 3: Mean per-joint rotation and position errors on AMASS and a real-world whole-body dataset with varying numbers of missing sensors. Loose-fitting models (Ours, LIP) are compared to tight-fitting ones (IMUPoser, PIP).
Based on the results presented in Table 2, we observe that models trained on simulated and synthetic data perform exceptionally well on the five AMASS datasets. However, their performance gradually declines on the LIP dataset, culminating in significantly reduced accuracy on the whole-body real-world dataset. In contrast, models trained on real-world data achieve good performance on the whole-body dataset but struggle with the AMASS and LIP datasets, likely due to the tighter fit conditions. This indicates that all diffusion models and both LIP models perform significantly better when the distribution of test data closely matches that of the training data, revealing a problem with adaptation to different garment conditions or body shapes.
In very challenging scenarios with noisy IMU observations, the models trained on real-world data (Figures 5 and 6) tend to converge towards the mean predictions (neutral pose with legs and arms in relaxed positions). This raises the question of whether the data augmentation approach and emphasis on realistic data are optimal, given that noise and ambiguity in the recorded signals not only mislead the model but also reduce motion dynamics.
We found that incorporating garment-related parameters while training the model on simulated loose data is an effective way to maintain the expressiveness of motion through less noisy observations and still account for potential uncertainties with the addition of these three values.
While we introduced height and BMI to determine whether a TallThin or ShortFat physique was used for the clothing simulation, it remains to be seen whether these parameters in real-world scenarios pertain more to garment-related characteristics or the participant’s body shape. Since we lacked detailed information about the participants’ physiques, we were unable to assess this aspect or evaluate each participant’s performance across a long sequence and varying garment styles. However, our experiments on shorter motion sequences, using the optimal selection of the three parameters, demonstrate an improved quantitative performance, as well as a clear qualitative advantage. This is illustrated in Figure 7. Both suitable and unsuitable garment-related information can impact pose estimation accuracy, potentially improving or degrading it, with joint angle errors fluctuating by ±5 degrees and joint position errors by up to ±10 cm (compared to the garment-unaware model). Finally, the selection of the three parameters began by choosing values within the training range—height between 160 and 180 cm, BMI between 22 and 30, and the $\gamma$ parameter taking one of six predefined values. When the parameters were assigned unrealistic real-world height and BMI values, the motion became more restricted and less dynamic, resembling the behavior of the model trained on real-world data with loosely fitting garments (Upper-body Real).
Experiments with missing sensors (Table 3) show that our model handles such conditions better than others. IMUPoser’s performance reflects conservative outputs with limited motion, yielding low error but reduced realism. It also exhibits noticeably more jitter, even when all sensors are available. PIP, on the other hand, performs well only when all sensors are present; its predictions degrade significantly with each additional missing sensor. All models exhibit increased jitter as the number of missing sensors grows [^1] Visual results are available at https://drive.google.com/ file/d/1QHEm2AaGohlv22clJdT9ElAMEb5aAtVc/view?usp=sharing.. The last two columns highlight the superior generalization ability of our models in real-world loose scenarios, where PIP and IMUPoser perform worse on Lorenz et al. Lorenz et al. (2022) dataset.
4.4 Ablation Study
Consistency regularization loss.
We attribute our model’s generalization ability and reduced jitter to the inclusion of this loss function. It compels the model to generate similar outputs for slightly varied and noisier versions of the observations, thereby enhancing its robustness. This is demonstrated by the degraded performance on real-world datasets and the enhanced performance on simulated datasets—which share the same characteristics as the training dataset—when the loss is omitted from the objective function. Table 4 presents the performance of the upper- and whole-body models trained on simulated data without this loss function.
Dataset   Metrics Upper-body Sim Whole-body Sim
AMASS dataset MPJRE 12.956 ± 5.674 12.693 ± 4.573
MPJPE 7.486 ± 5.255 8.130 ± 4.620
MPJVE 24.445 25.150
Jitter 130.772 106.773
GT Jitter 137.073 140.756
Real-world upper-body dataset MPJRE 21.701 ± 6.390 __
MPJPE 13.214 ± 7.645 __
MPJVE 43.738 __
Jitter 289.034 __
GT Jitter 33.412 __
Real-world whole-body dataset MPJRE 32.398 ± 7.362 25.758 ± 6.074
MPJPE 27.246 ± 15.289 28.647 ± 14.420
MPJVE 23.978 29.276
Jitter 50.890 72.931
GT Jitter 43.967 54.698
Table 4: Performance metrics of the conditional models trained on simulated data without $\mathcal{L}_{\textit{c}}$.
Importance of conditioning on all available signals.
We focused on conditional models that fully rely on provided IMU recordings, a dependency often perceived as their shortcoming. In contrast, an unconditional model can handle missing data, making it particularly useful in scenarios where unreliable signals are received from the root sensor and extremities. This approach allows for estimating ambiguous signals rather than conditioning predictions on noisy observations. Therefore, we trained an unconditional model to explore this potential improvement. During training, we adopted an inpainting method by masking the data from the root sensor, as the root joint estimations of the conditional models exhibited the largest angle errors. This approach is illustrated in Figure 8.
Figure 8: Unconditional inpainting-based diffusion model.
We generate a mask indicating for each element whether the loose-wear IMU observation is provided (0) or needs to be predicted by the model (1). Following the inpainting approach proposed by Cherel et al. Cherel et al. (2024), the model input consists of the masked ground truth data $x_{\text{masked}}$ (Equation 4), the noisy input, and the generated mask, concatenated along the sequence dimension. Temporal information $t$ is also included in the input to the denoiser. The model’s output is a comprehensive vector that includes pose estimates, loose-wear IMU data, tight-wear IMU data, and joint positions.
$\mathbf{x}_{\text{masked}}=(1-\text{mask})\mathbf{x}_{\text{0}}.$ (4)
The objective function mirrors that of the conditional whole-body model, with the addition of the L1 loss term, $\mathcal{L}_{\text{loose\_recon}}$, to account for errors in the reconstruction of the missing root sensor’s data. While including all observations led to similar performance in both the conditional and unconditional models, the results in Table 5 suggest that completely excluding the root sensor’s data from all frames during evaluation is not advisable, as this data is not consistently unreliable. The presence of reliable observations is likely a contributing factor to the better performance of the model when all data is available. By omitting certain data entirely, the model is granted the freedom to predict the pose of the affected joint by considering other sensors’ data that could be misleading itself (such as data obtained from the extremities), which can lead to inaccurate estimates, particularly pronounced when root rotation is minimal.
Metrics Conditional Unconditional
MPJRE 24.722 ± 5.701 25.933 ± 5.755
MPJPE 28.648 ± 14.47 31.021 ± 11.437
MPJVE 26.974 35.223
Jitter 34.800 150.606
GT Jitter 54.698 54.698
Root angle error 82.123 103.849
Table 5: Comparison between the conditional whole-body model (trained/evaluated on all six sensors) and an inpainting-based unconditional diffusion model (trained/evaluated with missing root sensor).
Pose, tightly attached IMU data, and joint positions in output vector.
We adopted the training scheme presented in Figure 4, which involves estimating not only the pose but also data from tightly attached sensors and joint positions. Training the model solely for pose estimation led to abrupt motions. Unlike the loss function in Equation 3, which includes six distinct losses, the pose-only model required additional loss components to reduce jitter. These included velocity, acceleration, and jerk losses over one, three, and five frames, as well as velocity losses for joint positions over the same intervals.
5 Conclusion
In this paper, we introduce a new task of full-body pose estimation from sparse and loosely attached IMU sensor data. Our approach builds on a multi-source training strategy that leverages simulated, synthetic, and real-world IMU data. We simulate IMU observations using garment-aware motion datasets and develop a diffusion model to synthesize realistic loose-wear IMU signals. This framework establishes a foundation for a comprehensive comparison of model performance across different data sources, allowing us to systematically evaluate the strengths and limitations of each. In addition, we demonstrate that incorporating garment-related parameters during training on simulated loose data enhances model expressiveness and helps address the inherent ambiguities of loosely attached IMU sensors. Experiments show our diffusion models outperform the state-of-the-art methods, providing a new direction for future research.